Applications of Differentiation

Consider the function f(x)= 2sin(x) + [(sqrt3)/2] x2 This function has two inflection numbers A<B in [0,2pi]:


A=    and   B=     For each of the following intervals, tell whether f(x) is concave up (type in CU) or concave down (type in CD).

[0,A)=

(A,B)=

(B,2pi]=
Answer #1

f(x)=cosx +(√2 /2)x

f '(x)=-sinx+(√2 /2)

for critical numbers f '(x)=0

-sinx+(√2 /2)=0

sinx=(√2 /2)

x=pi/4 ,3pi/4

A=pi/4

B=3pi/4

f "(x)=-cosx

f "(pi/4)=-cos(pi/4)= -(√2 /2) <0

f "(3pi/4)=-cos(3pi/4)= (√2 /2) >0

f(x) has a localmaximum at A and a local minimum at B


answered by: ANURANJAN SARSAM